# A Distributed Privacy Preserving Model for the Detection of Alzheimer’s Disease

### v3

@@ Abstract
BACKGROUND: Segmentation of medical data, concerns about personal health information (PHI) breaches, and the direct and indirect costs of consolidating and managing such segmented date should motivate diagnostic machine learning (DML) researchers to identify privacy-preserving machine learning algorithms that can train on distributed or decentralized datasets of different modalities. Federated learning models provide such a decentralized machine learning framework in which multiple investigators in possession of disparate datasets and working on different devices or servers can train collaboratively a global machine learning models without ever having to exchange local data and thus can meet statutory PHI protections. To this end, a vertical federate learning model is devised and tested for efficacy in the detection of Alzheimer’s Disease (AD).
METHODS: The second version of Open Access Series of Imaging Studies– with its panoply of demographic, imaging, and clinical assessment datasets– was used to test a multimodal vertical federated learning (VFL) model for AD detection.
RESULTS: By training and validating this VFL model on the demographic, clinical, and MRI data in OASIS-2, an 82.9% accuracy rate is achieved, consistent with previously reported results.
CONCLUSIONS: The VFL architecture proposed herein offers a novel distributed architecture, enabling collaborative learning across diverse sources of medical data while respecting statutory privacy constraints. By leveraging multiple modalities of data, the robustness and accuracy of AD detection can be enhanced. This model not only contributes to the advancement of federated learning techniques but also holds promise for overcoming the hurdles posed by data segmentation in medical research.
@@ 1 Introduction
The increasing volume of medical data should be a priceless asset for advancing diagnostic machine learning (DML) models in healthcare 1; 2; 3. Commonly, these large datasets are consolidated onto a centralized server for collaborators to investigate. Unfortunately, consolidating medical data in such a fashion in order to train a centralized diagnostic machine learning model can be burdensome, costly, and impractical. For example, the Health Insurance Portability and Accountability Act (HIPAA) 4; 5 and as updated through the Health Information Technology for Economic and Clinical Health Act (HITECH)6 mandate safeguards to protect any personal health information (PHI) from any breaches. Additionally, consolidating medical data in order to train a centralized model often can be an impractical endeavor. The segmentation inherent in medical data across diverse sources imposes significant limitations on both the depth and breadth of available information for robust DML training 7. Not surprisingly, consolidating and managing this data is fraught with administrative burdens, technical challenges, and considerable expenses. Specifically, regional health information organizations (RHIOs) require US \$12 million each for development and incur \$2-3 million in operating costs 8; 9. Just as a physician must grasp different types of clinical data (such as a history, physical examination, laboratory results, imaging studies, and questionnaires) to develop a unifying diagnosis, each data source used to train a DML model provides only a partial perspective. DML demands novel methodologies capable of training from different modes of data while optimizing model development in any distributed environment. Thus, the utility of multimodal models becomes clearly apparent.
@@
Arguably, as explained above, prevailing approaches to multimodal model training practice data consolidation 10 –fundamentally misaligned with the objectives of (1) preserving data segmentation, (2) protecting PHI, and (3) still optimizing model performance. On the other hand, the integration of diverse data modalities holds promise for a more comprehensive understanding, particularly in addressing complex medical conditions. Neuropsychiatric disorders, such as Alzheimer’s Disease (AD), require interpreting a host of different objective and subjective data to establish a definitive diagnosis. Because DML research on AD using multimodal data already has been reported, I chose to train a novel machine learning model– specifically, a vertical federated learning model– to detect AD. The decision to address AD not only stems from a technical standpoint but also aligns with a broader, urgent societal need. Specifically, in 2024, an estimated 6.9 million or 1 in 9 (11%) of all Americans aged 65 years and older are living with AD11. The Centers for Medicare and Medicaid Services shall pay out US\$ 206 billion for direct AD care12. Highlighting the overall the overall economic and associated burdens of AD to patients, their families, and society is the fact that there is an additionally estimated 18 billion unpaid hours for dementia-related care in 2022, valued at \$339.5 billion 13. The number of Americans aged 65 and older not only will increase exponentially in the United States but also elsewhere– as will the cost and other burdens of caring for AD. Early detection and intervention of AD should improve the care and cut the cost of care for AD. Therefore, there is a critical need for technical innovations facilitating early intervention and personalized care.
@@ 2.1 What Is Federated Learning?
Federated learning is a decentralized machine learning framework which allows multiple disparate parties– often on separate devices or servers– to collaboratively train a global machine learning model without ever having to exchange their local data 14. In scenarios wherever data security, patient privacy, and maintenance of data locality are necessary, federated learning can be a powerful tool. Rather than consolidating data in a central repository, federated learning leverages a federated architecture, where each data source (a ”client” or ”federate”) maintains its own dataset and participates in model training through local computations. The global model then is updated iteratively, aggregating the model updates generated by each client while preserving the privacy and security of individual data sources. Based on the framework architecture, federated learning models can be described as horizontal or vertical.
@@ 2.2 Horizontal Federated Learning
Horizontal Federated Learning (HFL) is defined as a scenario in which different parties share the same feature space but contain different data samples 15. After initialization of a global model, each client, possessing the same type of data, trains the model locally. Clients collaborate to improve the global model by sharing insights from their locally collected data with the central server. The central server then aggregates the models received from the clients, typically by averaging the model parameters. This averaging process ensures that the global model benefits from the diverse knowledge learned on different clients while preserving privacy. One common method used in this process is Federated Averaging (FedAvg) 14, where the central server orchestrates the federated learning process by specifically averaging the model updates and then redistributing the improved model to all clients. Therefore, HFL is effective in scenarios where multiple federates are collecting similar data samples from multiple federates (or entities) without having to consolidate this data onto a centralized server. In a clinical setting, one could imagine a scenario where there are three hospitals. These hospitals all contain the same fields of data (such as magnetic resonance imaging or MRI). Radiology Department A at one hospital may have MRIs for Patients 1-3 who were being evaluated for AD, Radiology Department B may MRIs of Patients 4-6, and Radiology Department C for Patients 7-9. In this case, an HFL model could be used for DML training since one mode of data (MRIs) from different patients are obtained from multiple federates in order to train a single neural network. Local proprietary interests of the data and PHI protections thereby would be preserved since the actual images and patient identification is neither stored centrally nor shared among the collaborators . The architecture of HFL is illustrated in figure 1.
@@ 2.3 Clinical Applications for Horizontal Federated Learning
There are many survey papers that both propose and implement HFL in the healthcare research 16; 17. In terms of primary research of HFL models for AD detection training, Khalil and colleagues 18 introduce a hardware accelerated horizontal federated multi-layer perceptron on field programmable gate arrays with 89% accuracy on blood plasma data from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) 19. They also used K-nearest neighbors (KNN), support vector machines, decision trees, and logistic regression analyses.
@@
Similarly, MRIs from the ADNI and Australian Imaging, Biomarkers and Lifestyle (AIBL) 20 datasets were utilized to implement a federated domain adaptation framework via Transformer (FedDAvT)21. This approach performed Unsupervised Deep Domain Adaptation and managed to achieve the following accuracy rates of: (1) 88.75% for the AD vs. Normal Controls (NC), (2) 69.51% for Mild Cognitive Impairment (MCI) vs. NC, and (3) 69.88% for AD vs. MCI binary classification tasks.
@@
A process called federated conditional mutual learning (FedCM)–which considers the similarity between clients– improved these DML models that trained on MRI imaging 22. Using MRIs from the first version of Open Access Series of Imaging Studies (OASIS 1) 23, ADNI, and AIBL datasets and then by extending upon FedAvg through FedCM, accuracy rates of 75.5% accuracy rate on OASIS-1 and of 76.0% on AIBL and ADNI were achieved.
@@
In each of the above studies of HFL models for AD detection, only one modality of data was used– whether blood plasma protein levels or MRIs. Nevertheless, some investigators have attempted to train HFLs on multimodal data. For example, Gao and colleagues performed a manifold projection of each electroencephalogram (EEG) signal, first by calculating a domain loss on each disparate signal and then by applying HFL to classify each signal24. Although only one type of testing was analyzed (EEGs), manifold projection is necessary since each EEG signal comes from a different area of the brain and thus can be considered multimodal data. Moreover, this paper confirmed that all user data privacy was preserved. Additionally, Peng and colleagues used federated graph convolution so that both imaging data and demographic data are converted into uniform high level features. A special graph generative adversarial network is then used to fill out missing nodes in the graph. Therefore, the multimodal data is projected onto a uniform feature space, where FedAVG is applied to the locally trained graph for detecting both Alzheimer’s Disease (AD) and Autism Spectrum Disorder (ASD) 25. Both of the technologies described above succeeded in their objectives, training either a single federated neural network using an HFL process which would not perform well without the manifold projection 24 or on multimodal data by using the same manifold projection25. Unfortunately, projecting features onto a different latent space inevitably results in loss of information. Additionally, by losing the structure of the data, it is not possible to apply algorithms that best suit the data (e.g. Convolutional Neural Networks on images). Finally, the manifold projection of certain datasets is not possible because it is extremely difficult to project images down to the same latent space as demographic features, for example.
@@ 2.4 Vertical Federated Learning
Vertical Federated Learning (VFL)– in contrast to HFL– addresses instances in which multiple federates share the feature space and each training example is spread across different entities 26. Clients possess different features of a common entity, and the goal is to train a model that leverages all features for improved prediction and analysis. This approach enables predictive modeling and analytics across heterogeneous, multimodal data sources while preserving data privacy and security 27.
@@ 2.5 Clinical Applications for Vertical Federated Learning
VFL for disease diagnosis is far less common. Nevertheless, two notable publications are worth discussion. Firstly, Zhang and colleagues 28 focused on the security of VFL systems using an example from the Federated Technology Enabler29 (FATE). Developed by Webank’s AI department, FATE remains the world’s first industrial-grade federated learning open source framework, enabling enterprises and institutions to collaborate on data while protecting data security and privacy. Ju et al. 30 adapted the manifold projection proposed in Gao et al. 24 and applied it to federated transfer learning allowing for increased diagnostic accuracy by leveraging data diversity.
@@ 3.1 Dataset and Preprocessing of the Data
The second version of Open Access Series of Imaging Studies (OASIS-2) is comprised of a longitudinal cohort of 150 individuals aged 60 to 96 years 31. All subjects were right-handed and consisted of both male and female subjects. Table 1 summarizes the list of features provided in the OASIS-2 dataset.
@@
SSE Socioeconomic status as assessed by the Hollingshead Index of Social Position 32
MMSE Mini-Mental State Examination score 33
CDR Clinical Dementia Rating 34
ASF Atlas scaling factor (unitless). Computed scaling factor that transforms native-space brain and skull to the atlas target (i.e., the determinant of the transform matrix) 35
eTIV Estimated total intracranial volume (cm3) 35
nWBV Normalized whole-brain volume, expressed as a percent of all voxels in the atlas-masked image that are labeled as gray or white matter by the automated tissue segmentation process 36
@@
MRI images were imported in the form of NIfTI (Neuroimaging Informatics Technology Imitative) images 37 and then converted into tensors with NiBabel 38. Although a FATE compatible data loader to load the entire NiFTI images could have been used and writing a dedicated dataloader for NiFTI images would have been more useful in a clinical setting, such a feat would require approximately 1 GB of data per batch since our batch size is 32 and each NIfTI image is approximately 30 MB. Therefore, in the interest of efficiency, the MRI images were pre-processed into 256x256 JPEG format for training purposes. The 70th slice, 100th slice, and 125th slice were used to form the 3 channels of our JPEG image. Cross sectional views of a representative MRI image from our dataset is shown in figure 4.
@@ 3.2 Experimental Model
Image data resided on Federate A. Demographic data along with the Clinical Dementia Rating resided on Federate B. On Federate A, the bottom model used a modified ResNet 18 architecture39, a deep CNN employing residual connections to prevent vanishing gradients, architecture where the last layer of ResNet 18 was connected to a dense layer with 10 outputs. The accuracy of 3D CNNs compared to 2D CNNs on MRIs has been reported previouisly.40 Because– on average– 2D CNNs outperformed 3D counterparts, 40 I opted to use a 2D CNN. On Federate B, the bottom model was a feed-forward neural network consisting of 12 inputs (each corresponding to a demographic feature), a hidden layer with 20 neurons, and an output layer with 10 neurons. These models were subsequently interconnected through an interactive layer with a size of 10 neurons. The top layer of this VFL model also was situated on Federate B because of the label’s residence. This top layer consisted of a feed-forward neural network with 10 inputs and the resultant output. The Framework’s architecture is illustrated in figure 5.
@@ 4 Results
Khalil et al. 18 ADNI Plasma 19 MLP HFL 89%
@@
Lei et al. 21 ADNI 19, AIBL 20 FedDAvT HFL 89%
Huang et al. 22 ADNI 19, OASIS 1 23, AIBL 20 FedCM HFL N/A
OASIS 1 23 FedCM HFL 74.5%
AIBL 20 FedCM HFL 76%
Mandal (2024) OASIS 2 31 ResNet + DNN VFL 82%

### v4

@@ Abstract
In the era of rapidly advancing medical technologies, the segmentation of medical data has become inevitable, necessitating the development of privacy preserving machine learning algorithms that can train on distributed data. Consolidating sensitive medical data is not always an option particularly due to the stringent privacy regulations imposed by the Health Insurance Portability and Accountability Act (HIPAA). In this paper, I introduce a HIPAA compliant framework that can train from distributed data. I then propose a multimodal vertical federated model for Alzheimer’s Disease (AD) detection, a serious neurodegenerative condition that can cause dementia, severely impairing brain function and hindering simple tasks, especially without preventative care. This vertical federated learning (VFL) model offers a distributed architecture that enables collaborative learning across diverse sources of medical data while respecting privacy constraints imposed by HIPAA. The VFL architecture proposed herein offers a novel distributed architecture, enabling collaborative learning across diverse sources of medical data while respecting statutory privacy constraints. By leveraging multiple modalities of data, the robustness and accuracy of AD detection can be enhanced. This model not only contributes to the advancement of federated learning techniques but also holds promise for overcoming the hurdles posed by data segmentation in medical research.
@@ 1 Introduction
The increasing volume of medical data should be a priceless asset for advancing diagnostic machine learning (DML) models in healthcare [1, 2, 3]. Commonly, these large datasets are consolidated onto a centralized server for collaborators to investigate. Unfortunately, consolidating medical data in such a fashion in order to train a centralized diagnostic machine learning model can be burdensome, costly, and impractical. For example, the Health Insurance Portability and Accountability Act (HIPAA) [4, 5] and as updated through the Health Information Technology for Economic and Clinical Health Act (HITECH) [6] mandate safeguards to protect any personal health information (PHI) from any breaches. Additionally, consolidating medical data in order to train a centralized model often can be an impractical endeavor. The segmentation inherent in medical data across diverse sources imposes significant limitations on both the depth and breadth of available information for robust DML training [7]. Not surprisingly, consolidating and managing this data is fraught with administrative burdens, technical challenges, and considerable expenses. Specifically, regional health information organizations (RHIOs) require US \$12 million each for development and incur \$2-3 million in operating costs [8, 9]. Just as a physician must grasp different types of clinical data (such as a history, physical examination, laboratory results, imaging studies, and questionnaires) to develop a unifying diagnosis, each data source used to train a DML model provides only a partial perspective. DML demands novel methodologies capable of training from different modes of data while optimizing model development in any distributed environment. Thus, the utility of multimodal models becomes clearly apparent.
@@
Arguably, as explained above, prevailing approaches to multimodal model training practice data consolidation [10] –fundamentally misaligned with the objectives of (1) preserving data segmentation, (2) protecting PHI, and (3) still optimizing model performance. On the other hand, the integration of diverse data modalities holds promise for a more comprehensive understanding, particularly in addressing complex medical conditions. Neuropsychiatric disorders, such as Alzheimer’s Disease (AD), require interpreting a host of different objective and subjective data to establish a definitive diagnosis. Because DML research on AD using multimodal data already has been reported, I chose to train a novel machine learning model– specifically, a vertical federated learning model– to detect AD. The decision to address AD not only stems from a technical standpoint but also aligns with a broader, urgent societal need. Specifically, in 2024, an estimated 6.9 million or 1 in 9 (11%) of all Americans aged 65 years and older are living with AD [11]. The Centers for Medicare and Medicaid Services shall pay out US\$ 206 billion for direct AD care [12]. Highlighting the overall the overall economic and associated burdens of AD to patients, their families, and society is the fact that there is an additionally estimated 18 billion unpaid hours for dementia-related care in 2022, valued at \$339.5 billion [13]. The number of Americans aged 65 and older not only will increase exponentially in the United States but also elsewhere– as will the cost and other burdens of caring for AD. Early detection and intervention of AD should improve the care and cut the cost of care for AD. Therefore, there is a critical need for technical innovations facilitating early intervention and personalized care.
@@ 2.1 What Is Federated Learning?
Federated learning is a decentralized machine learning framework which allows multiple disparate parties– often on separate devices or servers– to collaboratively train a global machine learning model without ever having to exchange their local data [14]. In scenarios wherever data security, patient privacy, and maintenance of data locality are necessary, federated learning can be a powerful tool. Rather than consolidating data in a central repository, federated learning leverages a federated architecture, where each data source (a ”client” or ”federate”) maintains its own dataset and participates in model training through local computations. The global model then is updated iteratively, aggregating the model updates generated by each client while preserving the privacy and security of individual data sources. Based on the framework architecture, federated learning models can be described as horizontal or vertical.
@@ 2.2 Horizontal Federated Learning
Horizontal Federated Learning (HFL) is defined as a scenario in which different parties share the same feature space but contain different data samples [15]. After initialization of a global model, each client, possessing the same type of data, trains the model locally. Clients collaborate to improve the global model by sharing insights from their locally collected data with the central server. The central server then aggregates the models received from the clients, typically by averaging the model parameters. This averaging process ensures that the global model benefits from the diverse knowledge learned on different clients while preserving privacy. One common method used in this process is Federated Averaging (FedAvg) [14], where the central server orchestrates the federated learning process by specifically averaging the model updates and then redistributing the improved model to all clients. Therefore, HFL is effective in scenarios where multiple federates are collecting similar data samples from multiple federates (or entities) without having to consolidate this data onto a centralized server. In a clinical setting, one could imagine a scenario where there are three hospitals. These hospitals all contain the same fields of data (such as magnetic resonance imaging or MRI). Radiology Department A at one hospital may have MRIs for Patients 1-3 who were being evaluated for AD, Radiology Department B may MRIs of Patients 4-6, and Radiology Department C for Patients 7-9. In this case, an HFL model could be used for DML training since one mode of data (MRIs) from different patients are obtained from multiple federates in order to train a single neural network. Local proprietary interests of the data and PHI protections thereby would be preserved since the actual images and patient identification is neither stored centrally nor shared among the collaborators . The architecture of HFL is illustrated in figure 1.
@@ 2.3 Clinical Applications for Horizontal Federated Learning
There are many survey papers that both propose and implement HFL in the healthcare research [16, 17]. In terms of primary research of HFL models for AD detection training, Khalil and colleagues [18] introduce a hardware accelerated horizontal federated multi-layer perceptron on field programmable gate arrays with 89% accuracy on blood plasma data from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) [19]. They also used K-nearest neighbors (KNN), support vector machines, decision trees, and logistic regression analyses.
@@
Similarly, MRIs from the ADNI and Australian Imaging, Biomarkers and Lifestyle (AIBL) [20] datasets were utilized to implement a federated domain adaptation framework via Transformer (FedDAvT) [21]. This approach performed Unsupervised Deep Domain Adaptation and managed to achieve the following accuracy rates of: (1) 88.75% for the AD vs. Normal Controls (NC), (2) 69.51% for Mild Cognitive Impairment (MCI) vs. NC, and (3) 69.88% for AD vs. MCI binary classification tasks.
@@
A process called federated conditional mutual learning (FedCM)–which considers the similarity between clients– improved these DML models that trained on MRI imaging [22]. Using MRIs from the first version of Open Access Series of Imaging Studies (OASIS 1) [23], ADNI, and AIBL datasets and then by extending upon FedAvg through FedCM, accuracy rates of 75.5% accuracy rate on OASIS-1 and of 76.0% on AIBL and ADNI were achieved.
@@
In each of the above studies of HFL models for AD detection, only one modality of data was used– whether blood plasma protein levels or MRIs. Nevertheless, some investigators have attempted to train HFLs on multimodal data. For example, Gao and colleagues performed a manifold projection of each electroencephalogram (EEG) signal, first by calculating a domain loss on each disparate signal and then by applying HFL to classify each signal [24]. Although only one type of testing was analyzed (EEGs), manifold projection is necessary since each EEG signal comes from a different area of the brain and thus can be considered multimodal data. Moreover, this paper confirmed that all user data privacy was preserved. Additionally, Peng and colleagues used federated graph convolution so that both imaging data and demographic data are converted into uniform high level features. A special graph generative adversarial network is then used to fill out missing nodes in the graph. Therefore, the multimodal data is projected onto a uniform feature space, where FedAVG is applied to the locally trained graph for detecting both Alzheimer’s Disease (AD) and Autism Spectrum Disorder (ASD) [25]. Both of the technologies described above succeeded in their objectives, training either a single federated neural network using an HFL process which would not perform well without the manifold projection [24] or on multimodal data by using the same manifold projection [25]. Unfortunately, projecting features onto a different latent space inevitably results in loss of information. Additionally, by losing the structure of the data, it is not possible to apply algorithms that best suit the data (e.g. Convolutional Neural Networks on images). Finally, the manifold projection of certain datasets is not possible because it is extremely difficult to project images down to the same latent space as demographic features, for example.
@@ 2.4 Vertical Federated Learning
Vertical Federated Learning (VFL)– in contrast to HFL– addresses instances in which multiple federates share the feature space and each training example is spread across different entities [26]. Clients possess different features of a common entity, and the goal is to train a model that leverages all features for improved prediction and analysis. This approach enables predictive modeling and analytics across heterogeneous, multimodal data sources while preserving data privacy and security [27].
@@ 2.5 Clinical Applications for Vertical Federated Learning
VFL for disease diagnosis is far less common. Nevertheless, two notable publications are worth discussion. Firstly, Zhang and colleagues [28] focused on the security of VFL systems using an example from the Federated Technology Enabler (FATE) [29]. Developed by Webank’s AI department, FATE remains the world’s first industrial-grade federated learning open source framework, enabling enterprises and institutions to collaborate on data while protecting data security and privacy. Ju et al. [30] adapted the manifold projection proposed in Gao et al. [24] and applied it to federated transfer learning allowing for increased diagnostic accuracy by leveraging data diversity.
@@ 3.1 Dataset and Preprocessing of the Data
The second version of Open Access Series of Imaging Studies (OASIS-2) is comprised of a longitudinal cohort of 150 individuals aged 60 to 96 years [31]. All subjects were right-handed and consisted of both male and female subjects. Table 1 summarizes the list of features provided in the OASIS-2 dataset.
@@
SSE Socioeconomic status as assessed by the Hollingshead Index of Social Position [32]
MMSE Mini-Mental State Examination score [33]
CDR Clinical Dementia Rating [34]
ASF Atlas scaling factor (unitless). Computed scaling factor that transforms native-space brain and skull to the atlas target (i.e., the determinant of the transform matrix) [35]
eTIV Estimated total intracranial volume (cm3) [35]
nWBV Normalized whole-brain volume, expressed as a percent of all voxels in the atlas-masked image that are labeled as gray or white matter by the automated tissue segmentation process [36]
@@
MRI images were imported in the form of NIfTI (Neuroimaging Informatics Technology Imitative) images [37] and then converted into tensors with NiBabel [38]. Although a FATE compatible data loader to load the entire NiFTI images could have been used and writing a dedicated dataloader for NiFTI images would have been more useful in a clinical setting, such a feat would require approximately 1 GB of data per batch since our batch size is 32 and each NIfTI image is approximately 30 MB. Therefore, in the interest of efficiency, the MRI images were pre-processed into 256x256 JPEG format for training purposes. The 70th slice, 100th slice, and 125th slice were used to form the 3 channels of our JPEG image. Cross sectional views of a representative MRI image from our dataset is shown in figure 4.
@@ 3.2 Experimental Model
Image data resided on Federate A. Demographic data along with the Clinical Dementia Rating resided on Federate B. On Federate A, the bottom model used a modified ResNet 18 architecture [39], a deep CNN employing residual connections to prevent vanishing gradients, architecture where the last layer of ResNet 18 was connected to a dense layer with 10 outputs. The accuracy of 3D CNNs compared to 2D CNNs on MRIs has been reported previouisly [40]. Because– on average– 2D CNNs outperformed 3D counterparts [40], I opted to use a 2D CNN. On Federate B, the bottom model was a feed-forward neural network consisting of 12 inputs (each corresponding to a demographic feature), a hidden layer with 20 neurons, and an output layer with 10 neurons. These models were subsequently interconnected through an interactive layer with a size of 10 neurons. The top layer of this VFL model also was situated on Federate B because of the label’s residence. This top layer consisted of a feed-forward neural network with 10 inputs and the resultant output. The Framework’s architecture is illustrated in figure 5.
@@ 4 Results
Khalil et al. [18] ADNI Plasma [19] MLP HFL 89%
@@
Lei et al. [21] ADNI [19], AIBL [20] FedDAvT HFL 89%
Huang et al. [22] ADNI [19], OASIS 1 [23], AIBL [20] FedCM HFL N/A
OASIS 1 [23] FedCM HFL 74.5%
AIBL [20] FedCM HFL 76%
Mandal (2024) OASIS 2 [31] ResNet + DNN VFL 82%